OpenStudy (anonymous):

Ok, I need a jumpstart. How do i solve this? 2[x+7]-7=1. This is absolute value. I keep getting the wrong answer.

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

you need to solve it until you only have the absolute value on one side

OpenStudy (anonymous):

I keep dividing by two. Im confused

OpenStudy (anonymous):

@Aysialove123 There is going to be two answers

OpenStudy (anonymous):

I'll show you the first couple steps @Starshine7

OpenStudy (anonymous):

Ok :) thanks, im ready

OpenStudy (anonymous):

Either (x+7)= 4 in which case x=-3 or -(x+7) = 4 in which case x= -11

OpenStudy (anonymous):

Original Problem\[2|x+7|-7=1\]Add 7 to each side\[2|x+7|=8\]Divide both sides by 2\[|x+7|=4\]What does this mean? It means that positive (x+7)=4 and negative (x+7)=4

OpenStudy (anonymous):

:0 oh!!! i remember now. because its absolute value i need to do it that way :3 thanks!!

OpenStudy (anonymous):

Let's do the positive first\[(x+7)=4\]Subtract 7\[x=-3\]

OpenStudy (anonymous):

\[-(x+7)=4\]\[-x-7=4\]\[-x=11\]\[x=-11\]

OpenStudy (anonymous):

You're Welcome

OpenStudy (anonymous):

:)